# Subgroups and diversity of left-orderable small cancellation groups

### 19 Dec 2023 v1

@@ 1. Introduction
A group is left-orderable if it admits a total order that is invariant under left-multiplication. Equivalently, a countable group is left-orderable if it acts by orientation preserving homeomorphisms on the real line. We are interested in left-orderability of negatively curved groups, such as hyperbolic groups or small cancellation groups. Many natural examples of such groups are left-orderable, for instance, non-abelian free groups, surface groups, one relator quotients of such groups [Bro80, Hem90, RW01, ADL11] and the fundamental groups of some hyperbolic $3$-manifolds [BRW05], see also [CR16]. Negatively curved left-orderable groups can also be produced as amalgamated free products or HNN-extensions of such groups under appropriate compatibility conditions [BG09, Theorems A and B], and as graphs of groups under the appropriate conditions [Chi11]. However, a ‘typical’ negatively curved group is not expected to be left-orderable. Indeed, random groups are not left-orderable in the density as well as the triangular model [Orl17, Orl21]. In addition, every small cancellation construction can easily be made not left-orderable by adding two relations to the presentation of the respective small cancellation group, see Example 3.3 below.
@@ 1.1. Subgroups of left-orderable small cancellation groups
Rips’ construction is a prominent tool to construct examples of negatively curved groups with various algebraic, geometric and algorithmic properties that contrast the properties of free or surface groups. This is often done by using the short exact sequence (see Theorem 1.1) to lift properties of the quotient $Q$ to $G$ or $N$. In view of Theorem 1.1, this method can be used to produce left-orderable groups with the respective properties. For example, there are appropriate left-orderable versions of [Rip82] (properties of subgroups and algorithmic properties), [BMS94], [BW01] (algorithmic and geometric properties), [GR20] and [Cou22, Theorem 1.3] (properties of growth functions and automorphisms). Thus, the properties of left-orderable small cancellation groups encompasses a wide range of properties of hyperbolic groups that are not satisfied by free or surface groups. In this sense, the left-orderable small cancellation groups, and the left-orderable hyperbolic groups, are rich classes of negatively curved groups.
@@ 1.3. Quasi-isometric diversity
Bowditch produced continuum many quasi-isometry classes of finitely generated small cancellation groups [Bow98]. Bowditch’s argument and [BG09, Theorem A] yields continuum many quasi-isometry classes of finitely generated left-orderable small cancellation groups, see Example 3.4 below. We combine the presentations of these groups with small cancellation presentations of left-orderable groups that are not locally indicable. This yields the following.
@@ 1.4. Perfect groups
Again, the groups in Proposition 1.9 can be realised as small cancellation groups. Thus, Proposition 1.9 implies Proposition 1.7. As mentioned in a personal communication by Arman Darbinyan, in a paper under preparation he is going to show quasi-isometric diversity of the class of left-orderable (in fact, bi-orderable) residually finite solvable groups.
@@
All groups constructed in this article split as an amalgamated product or an HNN-extension. It would be interesting to construct left-orderable small cancellation groups that do not split. Examples of small cancellation groups that do not split were given in [Pri83], see also [JW22, Lemma 5.1 and Remark 5.2]. These groups are not left-orderable. In fact, their presentations contain relators contradicting left-orderability, see Example 3.3 below. We also note that random groups in the density model do not split [DGP11].
@@ Acknowledgements
The author thanks Arman Darbinyan and Yash Lodha for valuable discussions and encouragement, and Yago Antolin, Goulnara Arzhantseva and Arman Darbinyan for useful comments on a previous version of this work. This work was supported by the Austrian Science Fund (FWF) project P 35079-N and the Mathematisches Forschungsinstitut Oberwolfach under its Research in Pairs program in year 2022 (Project ID 2212p).
@@ 2.1. Lexicographic orderings of the free group
Given a word $t$ as above, we proceed as follows. For $x$ and $y$ in $\mathcal{A}$ we set $x<_{t}y$ and $y^{-1}<_{t}x^{-1}$ if $x$ comes before $y$ in the word $t$ read from left to right. We also set $y<_{t}x^{-1}$ for all $x,y\in\mathcal{A}$. Next let $g$ be an element in the free group on $\mathcal{A}$ that is represented as a reduced word, i.e. in normal form. Let
@@ 3. Small cancellation groups
Next we discuss two examples of small cancellation groups. Our first example shows that every small cancellation construction can be made non-left-orderable.
@@
Let $G$ be a group generated by $x_{1},\ldots,x_{n}$. Let $r_{1}(x_{1},x_{2})$ be a positive word, i.e. the exponent of every letter $x_{1}$, $x_{2}$, is positive, and let $r_{2}(x_{1},x_{2})$ be a word such that every occurrence of $x_{1}$ has positive exponent, and every occurrence of $x_{2}$ has negative exponent. If $r_{1}$ and $r_{2}$ are relations in $G$, then $G$ is not left-orderable, as the subgroup generated by $x_{1}$ and $x_{2}$ is not left-orderable, see [CR16, Theorem 1.48]. Taking $r_{1}$ and $r_{2}$ appropriate small cancellation words, we observe that every small cancellation construction can be adapted to produce non-left-orderable groups by adding $r_{1}$ and $r_{2}$ to the presentation of the group.
@@ 4. Groups that are not locally indicable
This is analogous to the above arguments to prove that $\tau_{1}(gh_{i}g^{-1})=\tau_{2}(fk_{i}f^{-1})$.
@@ 5.2. Construction of left-orders
Case 2: $g=g_{0}x_{i}$, $p_{i}=x_{i}^{-1}$ and $gh_{i}^{-1}=g_{0}w_{i}^{-1}(a_{1}b_{1})$. Let us denote by $t$ the last letter of $g_{0}$ so that $g=g_{1}t^{\delta}x_{i}$. If $t<c$, we let $f=g_{0}^{\prime}x_{i}$. Then the assertion holds by definition. Otherwise $t\geq c$. If $t<x_{i}$ we let $f=g_{1}^{\prime}{t^{\prime}}^{\delta}c^{\delta}x_{i}$. Then the assertion holds by Lemma 2.2. Indeed, as $A<c\leq t$,
@@
$\tau(gx_{i}x^{\varepsilon})=\tau(g_{1}t^{\delta}x_{i}x^{\varepsilon})=\tau(g_{1}t^{\delta}c^{\delta}x_{i}x^{\varepsilon})=\tau^{\prime}(g_{1}^{\prime}{t^{\prime}}^{\delta}c^{\delta}x_{i}x^{\varepsilon})=\tau^{\prime}(f{x}^{\varepsilon}).$
@@
Case 4: $h_{i}=w(a_{1},a_{2})x_{i}^{\varepsilon}y^{\delta}x_{i}^{-\varepsilon}v(a_{1},a_{2})$ and $g=g_{0}u_{0}^{-1}(a_{1},a_{2})x_{i}^{\varepsilon}y^{-\delta}x_{i}^{-\varepsilon}w^{-1}(a_{1},a_{2})$, where $v=u_{0}u_{1}$ and $u_{1}$ is a non-trivial word. Then $gh_{i}=g_{0}u_{1}(a_{1},a_{2})$ and $k_{i}=\varphi(h_{i})=w(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{\delta}y_{i}^{-\varepsilon}v(b_{1},b_{2})$. In this case we let $f=g_{0}^{\prime}u_{0}^{-1}(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{-\delta}y_{i}^{-\varepsilon}w^{-1}(a_{1},b_{2})$.
@@
Next let $h=h_{i_{1}}\cdots h_{i_{l}}$, where either $l>1$ or $h_{i_{1}}\not=h_{i}$ and $h_{i_{1}}\not=h_{i}^{-1}$. If $l>1$, either $g$ is $p$-reduced with $h_{i_{1}}$ or $h_{i_{1}}=h_{i}$ and $gh_{i}$ is $p$-reduced with $h_{i_{2}}$. Similarly, either $g$ is $p$-reduced with $h_{i_{l}}^{-1}$ or $h_{i_{l}}=h_{i}$ and $gh_{i}$ is $p$-reduced with $h_{i_{l-1}}^{-1}$. Otherwise $g$ is $p$-reduced with $h_{i_{1}}$ and $h_{i_{1}}^{-1}$. The assertion from the Lemma thus follows from Lemma 5.4 given that the following hold for all pieces $q_{j}$ such that $h_{j}$ and $g$, or $h_{j}$ and $gh_{i}$, respectively, are $q_{j}$-reduced:
$\tau(gq_{j}a^{\varepsilon})=\tau^{\prime}(f\overline{q_{j}}a^{\prime\varepsilon})\hbox{ and }\tau(gx^{\varepsilon})=\tau^{\prime}(fx^{\varepsilon}),$
$\tau(gh_{i}q_{j}a^{\varepsilon})=\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\varepsilon})\hbox{ and }\tau(gh_{i}x^{\varepsilon})=\tau^{\prime}(fk_{i}x^{\varepsilon}).$
We prove that $\tau(gh_{i}q_{j}a^{\varepsilon})=\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\varepsilon})$. The proof of the other equalities is analogous and will not be given. We expand, using Lemma 2.3:
$\tau(gh_{i}q_{j}a^{\varepsilon})=\tau(g_{0}y^{\delta})+\tau(y^{\delta}x_{i}^{-\varepsilon})+\tau(x_{i}^{-\varepsilon}v(a_{1},a_{2})q_{j}a^{\varepsilon})$
@@
$\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\varepsilon})=\tau^{\prime}(f_{0}d_{2}^{\delta})+\tau^{\prime}(d_{2}^{\delta}y_{i}^{-\varepsilon})+\tau^{\prime}(y_{i}^{-\varepsilon}v(a_{1},a_{2})\overline{q_{j}}{a^{\prime}}^{\varepsilon}).$
@@
Case 4: $k_{i}=w(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{\delta}y_{i}^{-\varepsilon}v(b_{1},b_{2})$ and $g=g_{0}u_{0}^{-1}(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{-\delta}y_{i}^{-\varepsilon}w^{-1}(b_{1},b_{2})$, where $v=u_{0}u_{1}$ and $u_{1}$ is a non-trivial word. Then $gk_{i}=g_{0}u_{1}(b_{1},b_{2})$ and $h_{i}=\varphi^{-1}(k_{i})=w(a_{1},a_{2})x_{i}^{\varepsilon}y^{\delta}x_{i}^{-\varepsilon}v(a_{1},a_{2})$. In this case we let $f=g_{0}^{\prime}u_{0}^{-1}(a_{1},a_{2})x_{i}^{\varepsilon}y^{-\delta}x_{i}^{-\varepsilon}w^{-1}(a_{1},a_{2})$.
@@ 5.3. Embedding groups that are not locally indicable
As before, we let $z_{1}=a_{1}$, $z_{2}=a_{2}$, $z_{3}=b_{1}$, $z_{4}=b_{2}$, $z_{5}=c$, $z_{6}=d_{1}$, $z_{7}=d_{2}$, $z_{8}=d_{3}$, $z_{9}=e$, and let $z_{10}=e_{1}$, $z_{11}=e_{2}$, $z_{12}=e_{3}$ and $z_{13}=e_{4}$. Then we define the elements $h_{2i}$ and $k_{2i}$ by (2)-(8) and let $H$ be the subgroup freely generated by $h_{2},h_{4},\ldots,h_{2_{i}},\ldots$ and $K$ the subgroup freely generated by $k_{2},k_{4},\ldots,k_{2_{i}},\ldots$ respectively. As before we denote by $\varphi$ the natural isomorphism from $H$ to $K$ that sends $h_{2i}$ to $k_{2i}$, and let

### 29 Jan 2024 v2

@@ 1. Introduction
A group is left-orderable if it admits a total order that is invariant under left-multiplication. Equivalently, a countable group is left-orderable if it acts by orientation preserving homeomorphisms on the real line. Left-orderable groups attract a lot of attention in the context of dynamics and low dimensional topology, see e.g. [DNR14, CR16]. We are interested in left-orderability of negatively curved groups, such as hyperbolic groups or small cancellation groups. Many natural examples of such groups are left-orderable, for instance, non-abelian free groups, surface groups, one relator quotients of such groups [Bro80, Hem90, RW01, ADL11] and the fundamental groups of some hyperbolic $3$-manifolds [BRW05], see also [CR16]. Negatively curved left-orderable groups can also be produced as amalgamated free products or HNN-extensions of such groups under appropriate compatibility conditions [BG09, Theorems A and B], and as graphs of groups under the appropriate conditions [Chi11]. However, a ‘typical’ negatively curved group is not expected to be left-orderable. Indeed, random groups are torsion-free but not left-orderable in the density as well as the triangular model [Orl17, Orl21]. In addition, every torsion-free small cancellation construction can easily be made not left-orderable by adding two relations to the presentation of the respective group, see Example 3.3 below. Thus, there is a wealth of torsion-free finitely presented small cancellation groups that are not left-orderable, even though every finitely presented small cancellation group is commensurable to a subgroup of a right-angeled Artin group [Ago13, HW08, Wis04], hence, contains a finite index bi-orderable subgroup.
@@ 1.1. Subgroups of left-orderable small cancellation groups
Every finitely generated recursively presented group is a quotient by a finitely generated subgroup of a subgroup of a right-angled Artin group, hence, a bi-orderable group [AM15, Proposition 7.2]. However, in this construction the bi-orderable group is not necessarily a small cancellation or hyperbolic group, although it is a subgroup of a finitely presented small cancellation group, see [AM15, Proposition 7.2] and [HW08, Theorem 10.1].
Remark 1.4.
@@
Rips’ construction is a prominent tool to construct examples of negatively curved groups with various algebraic, geometric and algorithmic properties that contrast the properties of free or surface groups. This is often done by using the short exact sequence (see Theorem 1.1) to lift properties of the quotient $Q$ to $G$ or $N$. In view of Theorem 1.1, this method can be used to produce left-orderable small cancellation groups with the respective properties. For example, there are appropriate left-orderable versions of [Rip82] (properties of subgroups and algorithmic properties), [BMS94], [BW01] (algorithmic and geometric properties), [GR20] and [Cou22, Theorem 1.3] (properties of growth functions and automorphisms). Thus, the properties of left-orderable small cancellation groups encompass a wide range of properties of hyperbolic groups that are not satisfied by free or surface groups. In this sense, the left-orderable small cancellation groups, and the left-orderable hyperbolic groups, are rich classes of negatively curved groups.
@@ 1.3. Quasi-isometric diversity
Bowditch produced continuum many quasi-isometry classes of finitely generated small cancellation groups [Bow98]. Bowditch’s argument and [BG09, Theorem A] yield continuum many quasi-isometry classes of finitely generated left-orderable small cancellation groups, see Example 3.4 below. We combine the presentations of these groups with small cancellation presentations of left-orderable groups that are not locally indicable. This yields the following.
@@ 1.4. Perfect groups
The groups in Proposition 1.10 are realised as small cancellation groups, see Section 4. Thus, Proposition 1.10 implies Proposition 1.8. As mentioned in a personal communication by Arman Darbinyan, in a paper under preparation he is going to show quasi-isometric diversity of the class of left-orderable (in fact, bi-orderable) residually finite solvable groups.
@@
The results of this paper appear to be the first applications of small cancellation theory to the study of left-orderable groups. As all groups constructed in this article split as an amalgamated product or an HNN-extension, it would be interesting to construct left-orderable small cancellation groups that do not split. Examples of small cancellation groups that do not split were given in [Pri83], see also [JW22, Lemma 5.1 and Remark 5.2]. These groups are not left-orderable. In fact, their presentations contain relators contradicting left-orderability, see Example 3.3 below. We also note that random groups in the density model do not split [DGP11].
@@ Acknowledgements
The author thanks Arman Darbinyan and Yash Lodha for valuable discussions and encouragement, and Yago Antolin, Goulnara Arzhantseva, Arman Darbinyan, Thomas Delzant and Ashot Minasyan for useful comments on a previous version of this work. This work was supported by the Austrian Science Fund (FWF) project P 35079-N and the Mathematisches Forschungsinstitut Oberwolfach under its Research in Pairs program in year 2022 (Project ID 2212p).
@@ 2.1. Lexicographic orderings of the free group
Given a word $t$ as above, we proceed as follows. For $x$ and $y$ in $\mathcal{A}$ we set $x<_{t}y$ and $y^{-1}<_{t}x^{-1}$ if $x$ comes before $y$ in the word $t$ read from left to right. We also set $x^{-1}<_{t}y$ for all $x,y\in\mathcal{A}$. Next let $g$ be an element in the free group on $\mathcal{A}$ that is represented as a reduced word, i.e. in normal form. Let
@@ 3. Small cancellation groups
Next we discuss two examples of small cancellation groups. Our first example shows that every torsion-free small cancellation construction can be made non-left-orderable.
@@
Let $G$ be a group generated by $x_{1},\ldots,x_{n}$. Let $r_{1}(x_{1},x_{2})$ be a positive word, i.e. the exponent of every letter $x_{1}$, $x_{2}$, is positive, and let $r_{2}(x_{1},x_{2})$ be a word such that every occurrence of $x_{1}$ has positive exponent, and every occurrence of $x_{2}$ has negative exponent. In addition, we assume that $r_{1}$ and $r_{2}$ are no proper powers. If $r_{1}$ and $r_{2}$ are relations in $G$, then $G$ is not left-orderable, as the subgroup generated by $x_{1}$ and $x_{2}$ is not left-orderable, see [CR16, Theorem 1.48]. Taking $r_{1}$ and $r_{2}$ appropriate small cancellation words, we observe that every small cancellation construction can be adapted to produce non-left-orderable groups by adding $r_{1}$ and $r_{2}$ to the presentation of the group. As $r_{1}$ and $r_{2}$ are no proper powers, $r_{1}$ and $r_{2}$ do not produce any new torsion elements.
@@ 4. Groups that are not locally indicable
This is analogous to the above arguments to prove that $\tau_{1}(gh_{i}g^{-1})=\tau_{2}(fk_{i}f^{-1})$. Thus, by Lemma 4.6, $\tau_{1}(ghg^{-1})=\tau_{2}(f\varphi(h)f^{-1}).$
@@ 5.2. Construction of left-orders
Case 2: $g=g_{0}x_{i}$, $p_{i}=x_{i}^{-1}$ and $gh_{i}^{-1}=g_{0}w_{i}^{-1}(a_{1}b_{1})$. Then $p_{j}$ is trivial for all $j\not=i$. Let us denote by $t$ the last letter of $g_{0}$ so that $g=g_{1}t^{\delta}x_{i}$. If $t<c$, we let $f=g_{0}^{\prime}x_{i}$. Then the assertion holds by definition. Otherwise $t\geq c$. If $t<x_{i}$ we let $f=g_{1}^{\prime}{t^{\prime}}^{\delta}c^{\delta}x_{i}$. Then the assertion holds by Lemma 2.2. Indeed, as $A<c\leq t$,
@@
$\tau(gx^{\varepsilon})=\tau(g_{1}t^{\delta}x_{i}x^{\varepsilon})=\tau(g_{1}t^{\delta}c^{\delta}x_{i}x^{\varepsilon})=\tau^{\prime}(g_{1}^{\prime}{t^{\prime}}^{\delta}c^{\delta}x_{i}x^{\varepsilon})=\tau^{\prime}(f{x}^{\varepsilon}).$
Similarly, $\tau(ga^{\varepsilon})=\tau^{\prime}(fa^{\varepsilon})$, hence, $\tau(gp_{j}^{-1}a^{\varepsilon})=\tau^{\prime}(f{\overline{p_{j}}}^{-1}{a^{\prime}}^{\varepsilon})$, for all $j\not=i$.
@@
Case 4: $h_{i}=w(a_{1},a_{2})x_{i}^{\varepsilon}y^{\delta}x_{i}^{-\varepsilon}v(a_{1},a_{2})$ and $g=g_{0}u_{0}^{-1}(a_{1},a_{2})x_{i}^{\varepsilon}y^{-\delta}x_{i}^{-\varepsilon}w^{-1}(a_{1},a_{2})$, where $v=u_{0}u_{1}$ and $u_{1}$ is a non-trivial word that ends with the piece $p_{i}$. Then $gh_{i}=g_{0}u_{1}(a_{1},a_{2})$ and $k_{i}=\varphi(h_{i})=w(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{\delta}y_{i}^{-\varepsilon}v(b_{1},b_{2})$. In this case we let $f=g_{0}^{\prime}u_{0}^{-1}(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{-\delta}y_{i}^{-\varepsilon}w^{-1}(b_{1},b_{2})$.
@@
Next let $h=h_{i_{1}}\cdots h_{i_{l}}$, where either $l>1$ or $h_{i_{1}}\not=h_{i}$ and $h_{i_{1}}\not=h_{i}^{-1}$. If $l>1$, either $g$ is $p$-reduced with $h_{i_{1}}$ or $h_{i_{1}}=h_{i}$ and $gh_{i}$ is $p$-reduced with $h_{i_{2}}$. Similarly, either $g$ is $p$-reduced with $h_{i_{l}}^{-1}$ or $h_{i_{l}}=h_{i}$ and $gh_{i}$ is $p$-reduced with $h_{i_{l-1}}^{-1}$. Otherwise $g$ is $p$-reduced with $h_{i_{1}}$ and $h_{i_{1}}^{-1}$. The assertion of the lemma thus follows from Lemma 5.4 given that the following hold for all pieces $q_{j}$ such that $h_{j}$ and $g$, or $h_{j}$ and $gh_{i}$, respectively, are $q_{j}$-reduced:
$\tau(gq_{j}a^{\zeta})=\tau^{\prime}(f\overline{q_{j}}a^{\prime\zeta})\hbox{ and }\tau(gx^{\zeta})=\tau^{\prime}(fx^{\zeta}),$
$\tau(gh_{i}q_{j}a^{\zeta})=\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\zeta})\hbox{ and }\tau(gh_{i}x^{\zeta})=\tau^{\prime}(fk_{i}x^{\zeta}),$
for any choice of $\zeta\in\{\pm 1\}$. We prove that $\tau(gh_{i}q_{j}a^{\zeta})=\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\zeta})$. The proof of the other equalities is analogous and will not be given. We expand, using Lemma 2.3:
$\tau(gh_{i}q_{j}a^{\zeta})=\tau(g_{0}y^{\delta})+\tau(y^{\delta}x_{i}^{-\varepsilon})+\tau(x_{i}^{-\varepsilon}v(a_{1},a_{2})q_{j}a^{\zeta})$
@@
$\tau^{\prime}(fk_{i}\overline{q_{j}}a^{\prime\zeta})=\tau^{\prime}(f_{0}d_{2}^{\delta})+\tau^{\prime}(d_{2}^{\delta}y_{i}^{-\varepsilon})+\tau^{\prime}(y_{i}^{-\varepsilon}v(a_{1},a_{2})\overline{q_{j}}{a^{\prime}}^{\zeta}).$
@@
Case 4: $k_{i}=w(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{\delta}y_{i}^{-\varepsilon}v(b_{1},b_{2})$ and $g=g_{0}u_{0}^{-1}(b_{1},b_{2})y_{i}^{\varepsilon}d_{2}^{-\delta}y_{i}^{-\varepsilon}w^{-1}(b_{1},b_{2})$, where $v=u_{0}u_{1}$ and $u_{1}$ is a non-trivial word that ends with the piece $p_{i}$. Then $gk_{i}=g_{0}u_{1}(b_{1},b_{2})$ and $h_{i}=\varphi^{-1}(k_{i})=w(a_{1},a_{2})x_{i}^{\varepsilon}y^{\delta}x_{i}^{-\varepsilon}v(a_{1},a_{2})$. In this case we let $f=g_{0}^{\prime}u_{0}^{-1}(a_{1},a_{2})x_{i}^{\varepsilon}y^{-\delta}x_{i}^{-\varepsilon}w^{-1}(a_{1},a_{2})$.
@@ 5.3. Embedding groups that are not locally indicable
As before, we let $z_{1}=a_{1}$, $z_{2}=a_{2}$, $z_{3}=b_{1}$, $z_{4}=b_{2}$, $z_{5}=c$, $z_{6}=d_{1}$, $z_{7}=d_{2}$, $z_{8}=d_{3}$, $z_{9}=e$, and let $z_{10}=e_{1}$, $z_{11}=e_{2}$, $z_{12}=e_{3}$ and $z_{13}=e_{4}$. Then we define the elements $h_{2i}$ and $k_{2i}$ by (2)-(8). That is, in comparison with the previous section, we add words $h_{j_{i,l}}$ and $k_{j_{i,l}}$ of type (6) and (7) for $z_{l}=z_{l}^{\prime}=e_{1}$, $e_{2}$, $e_{3}$ or $e_{4}$ respectively. Then we let $H$ be the subgroup freely generated by $h_{2},h_{4},\ldots,h_{2_{i}},\ldots$ and $K$ the subgroup freely generated by $k_{2},k_{4},\ldots,k_{2_{i}},\ldots$ respectively. As before we denote by $\varphi$ the natural isomorphism from $H$ to $K$ that sends $h_{2i}$ to $k_{2i}$, and let